The density of gold is 19.30 g/cm³ at 20.0°C, and the coefficient of linear expansion is 14.3 x 10^-6 °C^-1. Compute the density of gold at 90°C.
1
Expert's answer
2020-11-16T06:37:23-0500

Let's assume that gols is an isotropic material. Then the volumetric thermal expansion coefficient is three times the linear coefficient:


"\alpha_V = 3\alpha_L = 3\cdot 14.3\times 10^{-6}\degree C^{-1} = 4.29\times 10^{-5}\degree C^{-1}"

Let's take a piece of gold of mass "m". Then its volume at "T_0 = 20.0\u00b0C" is:


"V_0 = \dfrac{m}{\rho_0}"

According to the ermal expansion law, the volume of mass 1 g at "T = 90\degree C" is:


"V = V_0 + \alpha_VV_0(T-T_0) = \dfrac{m}{\rho} \\"

where "\rho" is the new density. Obtain


"\rho = \dfrac{m}{V_0 + \alpha_VV_0(T-T_0)} = \dfrac{m}{\dfrac{m}{\rho_0} + \alpha_V\dfrac{m}{\rho_0}(T-T_0)}=\\
=\dfrac{\rho_0}{1 + \alpha_V(T-T_0)}"

Substituting the numbers, obtain:


"\rho = \dfrac{\rho_0}{1 + \alpha_V(T-T_0)} = \dfrac{19.3}{1 + 4.29\times 10^{-5}\cdot 70} \approx 19.24\space g/cm^3"

Answer. "19.24\space g/cm^3".